# publications.bib

### v3

@@ 2. Preliminaries on Shimura Varieties
In this section we will recall some facts about Shimura varieties which we will need later in this paper.
@@ 2.1. Shimura Varieties
We will mainly work with the following two types of Shimura varieties.
2.1.1. PEL type A and C
Let $(\mathcal{O}_{B},\ast,L,\langle\cdot,\cdot\rangle)$ be an integral PEL datum, where $B$ is a finite-dimensional semisimple $\mathbb{Q}$-algebra, $\ast$ is a $\mathbb{Q}$-linear involution of $B$, with fixed field $F$, $\mathcal{O}_{B}$ is a $\ast$-stable $\mathbb{Z}$-order of $B$, $L$ is a lattice with $\mathcal{O}_{B}$-actions, and $\langle\cdot,\cdot\rangle:L\times L\rightarrow\mathbb{Z}(1)$ is a non-degenerate alternating form such that $\langle bv,v^{\prime}\rangle=\langle v,b^{*}v^{\prime}\rangle$, for all $b\in\mathcal{O}_{B}$ and $v,v^{\prime}\in L$.
To our integral PEL datum, we have the following group scheme $\mathbf{G}$ over $\mathbb{Z}$ whose $R$-points, for each $\mathbb{Z}$-algebra $R$, are given by
$\mathbf{G}(R):=\{(g,r)\in\mathrm{End}_{\mathcal{O}_{B}\otimes_{\mathbb{Z}}R}(L\otimes_{\mathbb{Z}}R)\times R|\langle gv,gw\rangle=r\langle v,w\rangle\text{ for all }v,w\in L\otimes_{\mathbb{Z}}R\}.$
The PEL data is of type A if $(B\otimes_{F}\overline{F},\ast)$ is isomorphic to $\mathrm{End}(W)\times\mathrm{End}(W)^{\mathrm{op}}$ with $(a,b)^{\ast}=(b,a)$ for some vector space $W$. The PEL data is of type C if $(B\otimes_{F}\overline{F},\ast)$ is isomorphic to $\mathrm{End}(W)$ with $\ast$ being the adjoint map with respect to a symmetric bilinear form on $W$. We will assume from now on that we are in one of these cases.
We now further assume the data is unramified at $p$; namely, that each term in the decomposition $B_{\mathbb{Q}_{p}}=\prod_{\mathfrak{p}|p}B\otimes_{F}F_{\mathfrak{p}}$ is a matrix algebra over an unramified extension of $\mathbb{Q}_{p}$. We will thus moreover assume that $\mathcal{O}_{B}\otimes\mathbb{Z}_{p}$ is a maximal order in $B_{\mathbb{Q}_{p}}$, and $L$ is self-dual after localization at $p$. This can be arranged following [Lan, Remark 1.3.4.8]. Note that these conditions equivalently ensure that the group $G$ is unramified.
We will now briefly discuss what conditions we may need to impose on the prime $p$ so that the form of the local group $G$ satisfies the conditions in Table (1). Firstly, suppose we are in type A. Then the center $Z(B)=F_{c}$ is a quadratic imaginary extension of $F$. Let $n$ be the $\mathcal{O}_{B}$ rank of $L$. Observe that we will need to assume that the prime $p$ satisfies for all primes $\mathfrak{p}$ of $F$ above $p$,
@@ 2.1.1. PEL type A and C
These conditions imply that $G$ will be a similitude subgroup of $\prod_{\mathfrak{p}}G_{\mathfrak{p}}$ where $G_{\mathfrak{p}}$ is either $\mathrm{Res}_{F_{\mathfrak{p}}/\mathbb{Q}_{p}}(\mathbb{G}_{m}\times\mathrm{GL}_{n})$ or $\mathrm{GU}_{n}$ for an odd unitary group over $\mathbb{Q}_{p^{2}}$.
Now suppose we are in type C. Since the PEL data is unramified at $p$, we see that $B\otimes_{F}F_{\mathfrak{p}}$ is indefinite for all primes $\mathfrak{p}$ of $F$ above $p$, and thus $G$ will be a similitude subgroup of
$\prod_{\mathfrak{p}}\mathrm{Res}_{F_{\mathfrak{p}}/\mathbb{Q}_{p}}(\mathrm{GSp}_{2n}).$
Here, we will need to assume that the rank $n$ of $L$ as an $\mathcal{O}_{B}$ lattice is either $1$ or $2$ to satisfy the conditions in Table (1).
Both types of Shimura varieties will be moduli spaces of abelian varieties with extra structures, which we will briefly describe. Let $K^{p}\subset G(\mathbb{A}_{f}^{p})$ be an open compact subgroup. To any PEL data, the Shimura variety $\mathfrak{S}(\mathbf{G},X)_{K}$ over $\mathcal{O}_{E_{\mathfrak{p}}}$ is the scheme which represents the functor that associates to each locally Noetherian scheme $S$ over $\mathcal{O}_{E_{\mathfrak{p}}}$ the set of isomorphism classes of tuples $(A,\lambda,\iota,\eta^{p})$ consisting of
@@
    A section $\eta^{p}\in\Gamma(S,\mathrm{Isom}_{B}(L\otimes_{\mathbb{Z}}\mathbb{A}_{f}^{p},\hat{V}(A)^{p})/K^{p})$, where $\mathrm{Isom}_{B}(L\otimes_{\mathbb{Z}}\mathbb{A}_{f}^{p},\hat{V}(A)^{p})$ is the étale torsor of isomorphisms that maps $\langle$,$\rangle$ , ⟩ to a $\mathbb{A}_{f}^{p\times}$ multiple of the pairing on $\hat{V}(A)^{p}$ defined by the Weil pairing,
satisfying the Kottwitz determinant condition that $\mathrm{det}(b|\mathrm{Lie}(A))=\mathrm{det}(b|V^{-1,0})$ as polynomial functions on $\mathcal{O}_{B}$, where $V=L\otimes\mathbb{Q}$ and $V_{\mathbb{C}}=V^{-1,0}\oplus V^{0,-1}$ is the Hodge decomposition.
@@ 2.1.2. Compactifications
We will now recall some constructions from the theory of toroidal compactifications of PEL type Shimura varieties from [Lan]. To match the setting in [Lan], we will moreover assume from now on that the level structure $K$ is a principal congruence subgroup for some $N\geq 3$, namely
$K=K(N)=\{g\in\mathbf{G}(\hat{\mathbb{Z}})|g\equiv 1\pmod{N}\}.$
We first recall the definition of a split, symplectic and admissible filtration from [Lan, §5.2.1]. Let $R$ be a commutative ring.
@@
A split, symplectic and admissible filtration on $L\otimes_{\mathbb{Z}}R$ is a two-step filtration on $L\otimes_{\mathbb{Z}}R$ by $(\mathcal{O}_{B}\otimes_{\mathbb{Z}}R)$-submodules, i.e. we have
$0=Z_{-3}\subset Z_{-2}\subset Z_{-1}\subset L\otimes_{\mathbb{Z}}R,$
@@
    $Z_{-2}$ and $Z_{-1}$ are annihilators of each other under the pairing $\langle$,$\rangle$ , ⟩ induced from $L$.
Definition 2.2.
Let $M$ be a finite $B$-module. Since $B\simeq\prod_{i}B_{i}$ where each $B_{i}$ is simple, we have a decomposition $M\simeq M_{i}^{m_{i}}$ where $M_{i}$ is the unique simple left $B_{i}$-module. We call the tuple $(m_{i})$ the $B$-multi-rank of $M$.
Let $R=\hat{\mathbb{Z}}$ and suppose that we have a split symplectic admissible filtration $Z=Z_{\bullet}$ as above.
@@
We thus define a cusp label as a pair $(Z,\Phi)$, where $Z$ is a split symplectic admissible filtration on $L\otimes_{\mathbb{Z}}\hat{\mathbb{Z}}$, and $\Phi$ is a torus argument for $Z$. Note that this is the generalization of the cusp labels $(Z,X)$ considered in [CS2, §2.5.2], as for the PEL type A Shimura data they considered, the assumption of principal polarization means we can set $X=Y$, and the torus argument $\Phi$ is determined by the $\mathcal{O}_{F}$-isomorphism.
There is an action of $\mathbf{G}(\mathbb{A}_{f})$ on pairs $(Z,\Phi)$, as defined in [Lan, §5.4.3], and we define a cusp label at level $K$ to be a $K$-orbit of pairs $(Z,\Phi)$.
@@
To each cusp label $(Z,\Phi)$, we can associate a split torus $E_{\Phi}$ over $\mathbb{Z}$, as constructed by Lan in [Lan, §6.4]. Let $S_{\Phi}=\mathbb{X}^{*}(E_{\Phi})$. Let $S_{\Phi}^{\vee}:=\mathrm{Hom}_{\mathbb{Z}}(S_{\Phi},\mathbb{Z})$ be the $\mathbb{Z}$-dual of $S_{\Phi}$, and let $(S_{\Phi})^{\vee}_{\mathbb{R}}:=S^{\vee}_{\Phi}\otimes_{\mathbb{Z}}\mathbb{R}$. The $\mathbb{R}$-vector space $(S_{\Phi})^{\vee}_{\mathbb{R}}$ is isomorphic to the space of Hermitian pairings $|\cdot,\cdot|:(Y\otimes\mathbb{R})\times(Y\otimes\mathbb{R})\rightarrow\mathcal{O}_{B}\otimes\mathbb{R}$ by sending a Hermitian pairing |,|$\cdot$,$\cdot$|| ⋅ , ⋅ | to the function $y\otimes\phi(y^{\prime})\mapsto\mathrm{Tr}_{B/\mathbb{Q}}(|y,y^{\prime}|)$ in $\mathrm{Hom}_{\mathbb{Z}}(S_{\Phi},\mathbb{R})$ (c.f. [Lan, §6.2.5]).
@@
Thus, we have an $\mathbb{R}$-vector space $(S_{\Phi})_{\mathbb{R}}^{\vee}$ of Hermitian pairings, and we define $P_{\Phi}$ to be the subset of $(S_{\Phi})_{\mathbb{R}}^{\vee}$ corresponding to positive semi-definite Hermitian pairings with admissible radicals (see [Lan, Definition 6.2.5.4] and subsequent discussion for the precise definition of admissible radical). $P_{\Phi}$ will a rational polyhedral cone in $(S_{\Phi})_{\mathbb{R}}^{\vee}$. Moreover, to every torus argument $\Phi$ we can also associate a stabilizer group $\Gamma_{\Phi}$. We thus let $\Sigma_{\Phi}$ be a $\Gamma_{\Phi}$-admissible rational polyhedral cone decomposition of $P_{\Phi}$, as in [Lan, Definition 6.1.1.14].
@@
    A $\Gamma_{\Phi}$admissible smooth rational polyhedral cone decomposition $\Sigma_{\Phi}$ for each cusp $(Z,\Phi)$ so that the cone decompositions are pairwise compatible.
The precise definition and proof of existence of such smooth admissible rpcd is [Lan, §6.3.3.2,§6.6.3.3]. Associated to this admissible smooth rpcd, we have a toroidal compactification of $\mathfrak{S}(G,X)_{K}$, as in the following theorem of Lan [Lan, Theorem 6.4.1.1].

### v5

@@ 2. Preliminaries on Shimura Varieties
In this section, we will recall some facts about Shimura varieties and Igusa varieties which we will use to study the geometry of the Hodge-Tate period morphism in the next section. More specifically, in §2.1, we start with reviewing the definition of a Shimura datum of PEL type AC that we will be interested in (§2.1.1), and then we introduce the minimal and toroidal compactifications of the Shimura varieties attached to these datum (§2.1.3). In the next section §2.2, we carry out an analogous discussion for Igusa varieties, first reviewing their definition and moduli interpretation (§2.2.1), and then studying their toroidal and minimal compactifications (§2.2.4, §2.2.3). This will be important in our discussion in §3.1 for describing the fibers of the Hodge-Tate period morphism on the open Shimura variety.
@@ 2.1. Shimura Varieties
We will mainly work with the following two types of Shimura varieties. Our main reference here is [41, §1].
2.1.1. PEL type AC
Let $(\mathcal{O}_{B},\ast,L,\langle\cdot,\cdot\rangle,h)$ be an (integral) PEL datum, as defined in [41, §1.2.1]. Note here that by tensoring $O_{B}$ and $L$ by $\mathbb{Q}$, we obtain the (rational) PEL data as introduced by Kottwitz [39, §1-4]. More precisely, we require that $B$ is a finite-dimensional semisimple $\mathbb{Q}$-algebra, $\ast$ is a $\mathbb{Q}$-linear involution of $B$, with fixed field $F$, $\mathcal{O}_{B}$ is a $\ast$-stable $\mathbb{Z}$-order of $B$, $L$ is a lattice with $\mathcal{O}_{B}$-action, and $\langle\cdot,\cdot\rangle:L\times L\rightarrow\mathbb{Z}(1)$ is a non-degenerate alternating form such that $\langle bv,v^{\prime}\rangle=\langle v,b^{*}v^{\prime}\rangle$, for all $b\in\mathcal{O}_{B}$ and $v,v^{\prime}\in L$. Moreover $h$ is an $\mathbb{R}$-algebra homomorphism
$h:\mathbb{C}\rightarrow\mathrm{End}_{O_{B}\otimes_{\mathbb{Z}}\mathbb{R}}(L\otimes_{\mathbb{Z}}\mathbb{R}),$
such that
  1. (1)
    $\langle h(z)v,w\rangle=\langle v,h(\bar{z})w\rangle$ for all $v,w\in L\otimes_{\mathbb{Z}}\mathbb{R}$ and all $z\in\mathbb{C}$
  2. (2)
    the symmetric bilinear form $(v,h(i)w)$ on $L\otimes_{\mathbb{Z}}\mathbb{R}$ is positive definite.
To our integral PEL datum, we can associate the following group scheme $\mathbf{G}$ over $\mathbb{Z}$ whose $R$-points, for each $\mathbb{Z}$-algebra $R$, are given by
$\mathbf{G}(R):=\{(g,r)\in\mathrm{GL}_{\mathcal{O}_{B}\otimes_{\mathbb{Z}}R}(L\otimes_{\mathbb{Z}}R)\times R^{\times}|\langle gv,gw\rangle=r\langle v,w\rangle\text{ for all }v,w\in L\otimes_{\mathbb{Z}}R\}.$
We will assume from now on that $B$ only has simple factors of type AC. Here, we follow [41, Definition 1.2.1.15] classifying simple $\mathbb{Q}$-algebras with a positive involution. Observe that from [41, Proposition 1.2.3.11], this is equivalent to $\mathbf{G}^{\mathrm{ad}}$ having simple factors only of type A and C.
We now further assume the data is unramified at $p$; namely, that each term in the decomposition $B_{\mathbb{Q}_{p}}=\prod_{\mathfrak{p}|p}B\otimes_{F}F_{\mathfrak{p}}$ is a matrix algebra over an unramified extension of $\mathbb{Q}_{p}$. We will moreover assume that $\mathcal{O}_{B}\otimes\mathbb{Z}_{p}$ is a maximal order in $B_{\mathbb{Q}_{p}}$, and $L$ is self-dual after localization at $p$. This can be arranged following [41, Remark 1.3.4.8]. Note that these conditions equivalently ensure that the group $\mathbf{G}_{\mathbb{Z}_{p}}$ is a smooth reductive group scheme.
We will now briefly discuss what conditions we may need to impose on the prime $p$ so that the form of the local group $G:=\mathbf{G}_{\mathbb{Q}_{p}}$ satisfies the conditions in Table (1).
Type A: In this case, the center $Z(B)=F_{c}$ is a quadratic imaginary extension of $F$. Let $n$ be the $\mathcal{O}_{B}$ rank of $L$. Observe that we will need to assume that the prime $p$ satisfies for all primes $\mathfrak{p}$ of $F$ above $p$,
@@ 2.1.1. PEL type AC
These conditions imply that $G$ will be a similitude subgroup of $\prod_{\mathfrak{p}}G_{\mathfrak{p}}$ where $G_{\mathfrak{p}}$ is either $\mathrm{Res}_{F_{\mathfrak{p}}/\mathbb{Q}_{p}}(\mathbb{G}_{m}\times\mathrm{GL}_{n})$ or $\mathrm{GU}_{n}$ for either $n=2$ or an odd unitary group.
Type C: Since the PEL data is unramified at $p$, we see that $B\otimes_{F}F_{\mathfrak{p}}$ is indefinite for all primes $\mathfrak{p}$ of $F$ above $p$, and thus $G$ will be a similitude subgroup of
$\prod_{\mathfrak{p}}\mathrm{Res}_{F_{\mathfrak{p}}/\mathbb{Q}_{p}}(\mathrm{GSp}_{2n}),$
where here we define $n$ such that $2n$ is the $O_{B}$ rank of $L$, so $n$ is half the $O_{B}$ rank of $L$. In this case, we will need that $n$ is either $1$ or $2$ to satisfy the conditions in Table (1).
Finally, note that in the case of mixed type A and C, we can decompose $B=B_{1}\times B_{2}$, there $B_{1}$ is of type A, $B_{2}$ is of type C, and hence the $O_{B}$-lattice also decomposes as $L=L_{1}\oplus L_{2}$, where $L_{1}$ is an $O_{B_{1}}$-lattice, and $L_{2}$ is an $O_{B_{2}}$-lattice. We impose the relevant conditions on $n_{1}$, the $O_{B_{1}}$-rank of $L_{1}$, and $n_{2}$, half the $O_{B_{2}}$-rank of $L_{2}$. In this case, the group is a product of groups of the types listed above, and we let $n=n_{1}+n_{2}$.
To each integral PEL data we can associate a moduli space of abelian varieties with extra structures. Let $K^{p}\subset G(\mathbb{A}_{f}^{p})$ be an open compact subgroup, and let $K=K_{p}^{\mathrm{hs}}K^{p}$, where we recall $K_{p}^{\mathrm{hs}}\subset G(\mathbb{Q}_{p})$ denotes a hyperspecial level. To any PEL data, we let $\mathcal{M}_{K}$ over $\mathcal{O}_{E_{\mathfrak{p}}}$ be the scheme which represents the functor that associates to each scheme $S$ over $\mathcal{O}_{E_{\mathfrak{p}}}$ the set of isomorphism classes of tuples $(A,\lambda,\iota,\eta^{p})$ consisting of
@@
    A section $\eta^{p}\in\Gamma(S,\mathrm{Isom}_{B}(\underline{L\otimes_{\mathbb{Z}}\mathbb{A}_{f}^{p}},\underline{H^{1}(A,\mathbb{A}_{f}^{p}})/\underline{K^{p}}))$, where $\mathrm{Isom}_{B}(\underline{L\otimes_{\mathbb{Z}}\mathbb{A}_{f}^{p}},\underline{H^{1}(A,\mathbb{A}_{f}^{p}}))$ is the $\mathbf{G}(\mathbb{A}_{f}^{p})$ pro-étale torsor of $O_{B}\otimes\mathbb{A}_{f}^{p}$-equivariant isomorphisms that maps $\langle,\rangle$ to a $\underline{\mathbb{A}_{f}^{p\times}}$-multiple of the pairing on $H^{1}(A,\mathbb{A}_{f}^{p})$ defined by the Weil pairing.
satisfying the Kottwitz determinant condition that $\mathrm{det}(b|\mathrm{Lie}(A))=\mathrm{det}(b|V^{-1,0})$ as polynomial functions on $\mathcal{O}_{B}$, where $V=L\otimes\mathbb{Q}$ and $V_{\mathbb{C}}=V^{-1,0}\oplus V^{0,-1}$ is the Hodge decomposition. Here, the polynomial functions on $O_{B}$ for the action of $O_{B}$ on a finite locally free $O_{S}$-module $M$ for a scheme $S$ are defined in [39, §5].
Following [39, §8], we know that the generic fiber of $\mathcal{M}_{K}$ is a finite union of Shimura varieties for groups $\mathbf{G}^{\prime}$ with $\mathbf{G}(\mathbb{A})\simeq\mathbf{G}^{\prime}(\mathbb{A})$. In particular, the canonical integral model of the Shimura variety $\mathscr{S}(\mathbf{G},X)_{K}$ attached to $\mathbf{G}$ and the Hermitian space $X$ consisting of $\mathbf{G}(\mathbb{R})$-conjugates of $h$ is a union of connected components of $\mathcal{M}_{K}$. As such, we can still associate to $S$-points of $\mathscr{S}(\mathbf{G},X)_{K}$ an abelian scheme over $S$ with extra structures.
2.1.2. $p^{m}$-level
We will also consider normal integral models of Shimura varieties with deeper level at $p$, defined as follows. Assume from now on that the level structure $K$ is a principal congruence subgroup for some $N\geq 3$; namely, we have that
$K=K(N)=\{g\in\mathbf{G}(\hat{\mathbb{Z}})|g\equiv 1\pmod{N}\}$.
We consider Shimura varieties of level $K^{\prime}=K(p^{n}N)$ where $p\nmid N$. In this case, we consider $M_{K^{\prime}}$ the functor which assigns to every scheme $S$ over $E$ the set of isomorphism classes of tuples $(A,\lambda,\iota,\eta)$, where
  1. (1)
    AA is an abelian scheme over $S$, $\lambda$ a prime-to-$p$ polarization
  2. (2)
    $\iota:B\hookrightarrow\mathrm{End}(A)\otimes\mathbb{Q}$
  3. (3)
    $\eta$ is a section of the pro-étale torsor $\Gamma(S,\mathrm{Isom}_{B}(\underline{L\otimes_{\mathbb{Z}}\mathbb{A}_{f}},\underline{H^{1}(A,\mathbb{A}_{f}})/\underline{K^{\prime}}))$.
As in the case of the integral model at hyperspecial level, this moduli problem is representable for $N>3$ by a scheme over $E$. We can thus consider in this situation the composition
$M_{K^{\prime}}\rightarrow M_{K}\rightarrow\mathcal{M}_{K},$
and following [42] we let $\mathcal{M}_{K^{\prime}}$ denote the normalization of $\mathcal{M}_{K}$ in $M_{K^{\prime}}$, which by definition is a flat normal scheme over $\mathcal{O}_{E_{\mathfrak{p}}}$. Similar to before, $\mathcal{M}_{K^{\prime}}$ is a normal integral model for a finite union of Shimura varieties $\mathrm{Sh}_{K^{\prime}}(\mathbf{G},X)$ for groups $\mathbf{G}^{\prime}$ with $\mathbf{G}(\mathbb{A})\simeq\mathbf{G}^{\prime}(\mathbb{A})$, and we thus obtain a normal integral model $\mathscr{S}(\mathbf{G},X)_{K^{\prime}}$ for $\mathrm{Sh}_{K^{\prime}}(\mathbf{G},X)$.
@@ 2.1.3. Compactifications
We will now recall some constructions from the theory of toroidal compactifications of PEL type Shimura varieties from [41, 42].
We first recall the definition of a split, symplectic and admissible filtration from [41, §5.2.1]. Let $R$ be a commutative ring.
@@
A split, symplectic and admissible filtration on $L\otimes_{\mathbb{Z}}R$ is a two-step filtration on $L\otimes_{\mathbb{Z}}R$ by $(\mathcal{O}_{B}\otimes_{\mathbb{Z}}R)$-submodules, i.e. there is a flag
$0=Z_{-3}\subset Z_{-2}\subset Z_{-1}\subset L\otimes_{\mathbb{Z}}R=Z_{0},$
@@
    $Z_{-2}$ and $Z_{-1}$ are annihilators of each other under the pairing $\langle,\rangle$ induced from $L$.
Let $R=\hat{\mathbb{Z}}$ and suppose that we have a split symplectic admissible filtration $Z=Z_{\bullet}$ as above. The following is [41, Definition 5.4.1.3]:
@@
In the above definition, we use the notion of $B$-multi-rank as defined in [41, Definition 1.2.1.20].
Definition 2.3.
A cusp label is a pair $(Z,\Phi)$, where $Z$ is a split symplectic admissible filtration on $L\otimes_{\mathbb{Z}}\hat{\mathbb{Z}}$, and $\Phi$ is a torus argument for $Z$.
There is an action of $\mathbf{G}(\mathbb{A}_{f})$ on the set of possible cusp labels, defined as follows. We can consider rational cusp labels $(Z\otimes\mathbb{Q},\Phi\otimes\mathbb{Q})$, where $Z\otimes\mathbb{Q}$ is the filtration on $L\otimes\mathbb{A}_{f}$ obtained by tensoring with $\mathbb{Q}$ over $\mathbb{Z}$, and $\Phi\otimes\mathbb{Q}$ is similarly the tuple obtained by tensoring $\varphi_{-2},\varphi_{0}$ with $\mathbb{Q}$. Observe that this association with rational cusp labels is a bijection, since we can recover the integral filtration by intersecting with $L\otimes_{\mathbb{Z}}\hat{\mathbb{Z}}$ and the integral isomorphisms $\varphi_{-2}$ and $\varphi_{0}$ by restriction, since $\mathrm{Gr}_{-i}^{Z}$ is a lattice in $\mathrm{Gr}_{-i}^{Z\otimes\mathbb{Q}}$. Every $g\in\mathbf{G}(\mathbb{A}_{f})$ defines an automorphism of $L\otimes\mathbb{A}_{f}$, so given a filtration $Z\otimes\mathbb{Q}$, we let $g(Z\otimes\mathbb{Q})$ be the filtration obtained as the image under $g$, and $g(\Phi\otimes\mathbb{Q})$ the torus argument obtained by precomposing $\varphi_{-2},\varphi_{0}$ with the isomorphisms $\mathrm{Gr}_{-i}^{g(Z\otimes\mathbb{Q})}\simeq\mathrm{Gr}_{-i}^{Z\otimes\mathbb{Q}}$ induced by $g$.
We define a cusp label of level $K(N)$ to be a $K(N)$-orbit of cusp labels.
Remark 2.4.
@@
Note that this is the generalization of the cusp labels $(Z,X)$ considered in [8, §2.5.2], as for the PEL type A Shimura data they considered, the assumption of principal polarization means we can set $X=Y$, and the torus argument $\Phi$ is determined by the $\mathcal{O}_{F}$-isomorphism in part (2) of loc. cit. This definition is not the same as [41, Definition 5.4.1.9] for level $N$ (Lan uses $n$ as notation for the level). However, if we take a $K(N)$-orbit of cusps labels as defined here, we can look at the reduction mod $N$ of $\varphi_{-2}$ and $\varphi_{0}$ for any element in the orbit, and similarly taking the filtration mod $N$ gives a filtration on $L/NL$, and this recovers Lan’s definition.
@@
To each cusp label $(Z,\Phi)$, we can associate a split torus $E_{\Phi}$ over $\mathbb{Z}$, as constructed by Lan in [41, §6.2.2]. Let $S_{\Phi}=\mathbb{X}^{*}(E_{\Phi})$, and note that from the construction it is a quotient of $Y\otimes_{\mathbb{Z}}X$. Let $S_{\Phi}^{\vee}:=\mathrm{Hom}_{\mathbb{Z}}(S_{\Phi},\mathbb{Z})$ be the $\mathbb{Z}$-dual of $S_{\Phi}$, and let $(S_{\Phi})^{\vee}_{\mathbb{R}}:=S^{\vee}_{\Phi}\otimes_{\mathbb{Z}}\mathbb{R}$. The $\mathbb{R}$-vector space $(S_{\Phi})^{\vee}_{\mathbb{R}}$ is isomorphic to the space of Hermitian pairings $|\cdot,\cdot|:(Y\otimes\mathbb{R})\times(Y\otimes\mathbb{R})\rightarrow\mathcal{O}_{B}\otimes\mathbb{R}$ with isomorphism constructed by sending a Hermitian pairing $|\cdot,\cdot|$ to the function $y\otimes\phi(y^{\prime})\mapsto\mathrm{Tr}_{B/\mathbb{Q}}(|y,y^{\prime}|)$ in $\mathrm{Hom}_{\mathbb{Z}}(S_{\Phi},\mathbb{R})$ for any $y,y^{\prime}\in Y$ (c.f. [41, §6.2.5]).
@@
Thus, we have an $\mathbb{R}$-vector space $(S_{\Phi})_{\mathbb{R}}^{\vee}$ of Hermitian pairings, and we define $P_{\Phi}$ to be the subset of $(S_{\Phi})_{\mathbb{R}}^{\vee}$ corresponding to positive semi-definite Hermitian pairings with admissible radicals (see [41, Definition 6.2.5.4] and subsequent discussion for the precise definition of admissible radical). $P_{\Phi}$ will be a rational polyhedral cone in $(S_{\Phi})_{\mathbb{R}}^{\vee}$. Moreover, to every cusp $(Z,\Phi)$ we can also associate a stabilizer group $\Gamma_{Z}$ defined as follows. $\Gamma_{Z}\subset\mathrm{GL}_{O_{B}}(X)\times\mathrm{GL}_{O_{B}}(Y)$ is the group of pairs of isomorphisms $\gamma_{X}$ and $\gamma_{Y}$ of $X$ and $Y$ respectively satisfying $\phi=\gamma_{X}\circ\phi\circ\gamma_{Y}$, and mod $N$ we have $\varphi_{-2,N}=\gamma_{X,N}^{\vee}\circ\varphi_{-2,N}$ and $\varphi_{0,N}=\gamma_{Y,N}\circ\varphi_{0,N}$. Finally, let $\Sigma_{Z}$ be a $\Gamma_{Z}$-admissible rational polyhedral cone decomposition of $P_{Z}$, as in [41, Definition 6.1.1.14].
@@
    A $\Gamma_{Z}$ -admissible smooth rational polyhedral cone decomposition $\Sigma_{\Phi}$ for each cusp $(Z,\Phi)$ so that the cone decompositions are pairwise compatible.
The precise definition and proof of existence of such smooth admissible rpcd is [41, §6.3.3.2,§6.6.3.3]. Associated to this admissible smooth rpcd, we have a toroidal compactification of $\mathscr{S}(G,X)_{K}$ for $K=K(N)$ as constructed by Lan [41, Theorem 6.4.1.1] (when $p\nmid N$), and [42, §7-10] (when $p\mid N$).
The toroidal compactification may be described as follows. To a cusp label $(Z,\Phi)$ of level $K$, there is a smaller-dimensional Shimura variety, denoted $\mathscr{S}_{Z}$, which we will sketch briefly below, and whose precise definition may be found in [41, §6.2] $(p\nmid N)$ and [42, §8] ($p\mid N$). Then, we have the following: